OpenStudy (emilyjones284):

Im a bit confused on this one...ill give a medal (:

OpenStudy (emilyjones284):

\[2x=\sqrt{20+2x}\]

OpenStudy (anonymous):

x = 5/2

OpenStudy (emilyjones284):

thanks! would you mind showing me how you got that?

OpenStudy (anonymous):

square both sides and get \[4x^2=20+2x\] then solve the quadratic equation

OpenStudy (anonymous):

Back sorry, like satelite said square both sides,. By squaring both side it allows you to get rid of the square root

OpenStudy (anonymous):

Before you squaring the two sides you have to be careful that : \[2x\geq0\text{ et } 20+2x\geq 0\]

OpenStudy (emilyjones284):

thank you!